# Video Understanding with Large Language Models: A Survey

### v7

@@ II-B Background for LLMs
  • Scaling Laws [151]: With a significant extension of the model size (number of parameters), the pertaining data size, and computational resources, the performance of the model exhibits a pattern of regular growth, which can help researchers and engineers predict performance improvements or make effective decisions on model design and training.
  • Emergent Abilities [77]: When the parameter size and volume of training data for a large language model exceed a certain magnitude, some novel capabilities emerge, such as in-context learning, instruction following, and step-by-step reasoning. In-context learning allows the model to learn and make predictions based on the context provided within the input text without requiring explicit retraining. Instruction following enables the model to perform tasks based on natural language instructions. Step-by-step reasoning, e.g. chain-of-thought (CoT), allows the model to follow a logical sequence of steps to arrive at a conclusion, which is particularly useful for solving complex problems.
LLMs possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodal Large Language Models (MLLMs) [152, 76, 153, 154, 155] typically incorporate multimodal encoders, cross-modality aligners, and an LLM core structure. By combining multimodal encoders with the LLM, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Summarizer: In this subcategory, the primary function of the LLM is to summarize the analysis obtained from the Video Analyzers. The summarization approach varies based on the prompts provided to the LLM, ranging from highly condensed summary texts and captions to comprehensive summaries for answering specific questions. Notably, in Video Analyzer $\times$ LLM as Summarizer systems, the information flow is usually unidirectional (see Figure 4), with data flowing from the video to the Video Analyzer and then to the LLM, without any reverse process. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Summarizer category include: LaViLa [168], VLog [167], VAST [105], AntGPT [166], VIDOSC [165], Grounding-Prompter [164], LLoVi [163], Video ReCap [162], MVU [161], LangRepo [160], IG-VLM [159], MoReVQA [158], MM-Screenplayer [157], GIT-LLaVA [156], etc.
  • LLM as Manager: In this subcategory, the LLM primarily coordinates the overall system’s operation. It may actively generate commands to invoke various Video Analyzers to produce the desired results for the user, call the Video Analyzer and further process the obtained analysis before returning it to the user, or engage in multiple rounds of interaction with the Video Analyzer. Compared to the LLM as Summarizer category, the LLM as Manager category is more flexible and can be distinguished by its information flow complexity. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Manager category include: ViperGPT [190], Video ChatCaptioner [278], ChatVideo [189], AssistGPT [188], HuggingGPT [178], Hawk [185], ProViQ [187], LifelongMemory [186], DoraemonGPT [184], KEPP [183], VURF [182], VideoAgent (by Stanford) [181], VideoAgent (by PKU) [180], TV-TREES [179], SCHEMA [177], RAVA [176], GPTSee [172], TraveLER [175], LAVAD [174], VideoTree [173], LVNet [171], OmAgent [170], DrVideo [169], etc.
@@ III-A2 Video Embedder $\times$ LLM
A Video Embedder typically refers to a visual backbone/video encoder, such as ViT or CLIP, used to convert input videos into vector representations, known as video embeddings or video tokens. Some Embedders encode other modalities within the video, such as audio (e.g., CLAP [279]), which are also categorized under Video Embedder here (note that we do not consider the LLM’s tokenizer as an embedder). Unlike the text generated by the Video Analyzer, the vectors generated by the Video Embedder cannot be directly utilized by the LLM and usually require an adapter to map these embeddings from the vision’s (or other modalities’) semantic space to the text semantic space of the LLM’s input tokens.
@@
  • LLM as Text Decoder: In this subcategory, the LLM receives embeddings from the Video Embedder as input and decodes them into text outputs based on prompts or instructions. These tasks generally do not require fine-grained understanding or precise spatiotemporal localization, focusing mainly on general QA or captioning. Thus, the Vid-LLM behaves like a standard LLM during decoding. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Text Decoder category include: VideoLLM [223], Otter [247], Video-LLaMA [245], Video-ChatGPT [99], VALLEY [244], Macaw-LLM [243], MovieChat [241], Video-LLaVA [239], Chat-UniVi [237], Vista-LLaMA [236], VILA [235], GPT4Video [265], MovieLLM [225], InternVideo2 [224], MiniGPT4-Video [215], VideoChat2 [203], VideoLLaMA 2 [201], etc. See Figure 5 for the complete list.
  • LLM as Regressor: In this subcategory, the LLM receives embeddings from the Video Embedder as input and, like the Text Decoder, can output text. However, unlike the Text Decoder, the LLM as Regressor can also predict continuous values, such as timestamp localization in videos and bounding box coordinates for object trajectories, functioning similarly to a regressor performing regression tasks, even though it is fundamentally performing classification. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Regressor category include: VTimeLLM [258], SeViLA [255], TimeChat [257], GroundingGPT [256], OmniViD [254], LITA [253], HawkEye [252], Elysium [251], AVicuna [120], V2Xum-LLaMA [100], VLM4HOI [250], VideoLLM-online [249], Holmes-VAD [248], etc.
  • LLM as Hidden Layer: In this subcategory, the LLM also receives video embeddings as input but does not directly output text. Instead, it connects to a specially designed task-specific head to perform actual regression tasks, such as event time localization or object bounding box prediction in videos, while maintaining the LLM’s text output capability. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Hidden Layer category include: GPT4Video [265], OneLLM [264], VidDetours [263], Momentor [262], VTG-GPT [261], VITRON [260], VTG-LLM [259], etc.
@@
This category of Vid-LLMs is relatively rare. As the name suggests, it involves simultaneously using a Video Analyzer to obtain textual analysis of the video and a Video Embedder to encode the video into embeddings. The LLM receives both types of inputs along with other prompts/instructions and outputs responses to complete tasks. The subcategories here can flexibly be any of the Summarizer/Manager/Text Decoder/Regressor/Hidden Layer categories. Vid-LLMs in the (Analyzer + Embedder) $\times$ LLM category include: Vid2Seq [271], VideoChat [270], MM-VID [266], Auto-AD II [275], Vamos [274], PG-Video-LLaVA [277], MM-Narrator [273], SUM-shot [267], Merlin [269], Uni-AD [272], Vriptor [268], etc.
@@ III-B1 Training-free Vid-LLMs
Many Vid-LLMs systems are built on powerful LLMs with strong zero-shot, in-context learning, and Chain-of-Thought capabilities. These systems do not require training in the parameters of the LLM or other modules. Most Vid-LLMs in the Video Analyzer $\times$ LLM category are training-free because the information from the video and other accompanying modalities has already been parsed into text. At this point, the video understanding task has been transformed into a text understanding task. Since LLMs can unify almost all NLP tasks into generation tasks, they can also handle many video understanding tasks. SlowFast-LLaVA [280] is a training-free Video LLM that uses a two-stream input design to capture both spatial semantics and temporal context without fine-tuning and demonstrates capabilities across various video understanding benchmarks.
@@
Socratic Models [281] Varying CLIP ViT-L/14 - 1 V100 GPU RoBERTa, GPT-3 - 05/2022
LaViLa [168] 4 TimeSformer-B/L Cross-Attention 32 V100 GPUs GPT-2 XL frozeen - 12/2022
Vid2Seq [271] 100 CLIP ViT-L/14 Transformer Encoder 64 TPU v4 T5 0.2B 02/2023
ViperGPT [190] - - - - - - 03/2023
Vid ChatCaptioner [278] 100 - BLIP-2 24G GPU ChatGPT 20B 04/2023
VLog [167] - CLIP ViT-G - - ChatGPT 20B 04/2023
ChatVideo [189] - - - - ChatGPT 20B 04/2023
VideoChat [270] 4-32 ViT-G MLP+Q-fomer 1 A10 GPU StableVicuna 7B 05/2023
ChatBridge [246] 4 ViT-G Perceiver 8 A100 GPUs GPT-4 - 05/2023
VAST [105] 4,8 EVA-CLIP, BEATs, BERT-B - 64 V100 GPUs Vicuna 13B 05/2023
Otter [247] 4-8 CLIP ViT-L/14 - 4 RTX-3090 GPUs LLaMA 7B 05/2023
@@
VideoLLM [223] Varying 7 Task-Specific Video Encoders Linear Layer - GPT-2/T5/OPT/LLaMA 1.5/6.5/6.7/7B 05/2023
AssistGPT [188] 1/3 FPS - - 4 A5000 GPUs Vicuna 7B 06/2023
Video-LLaMA [245] 8 CLIP ViT-G Video Q-former - Vicuna 7B 06/2023
@@
Video-ChatGPT [99] 100 CLIP ViT-L/14 Linear Layer 8 A100 GPUs Vicuna-v1.1 7B 06/2023
LLMVA-GEBC [242] 96 CLIP ViT-G Video Q-former 2 A6000 GPUs OPT 13B 06/2023
mPLUG-video [96] 8 TimeSformer - - GPT/Blood - 06/2023
VALLEY [244] Varying CLIP ViT-L/14 Projection Layer 8 A100 GPUs StableVicuna 7B/13B 06/2023
@@
Macaw-LLM [243] Varying CLIP VIT-B/16 Alignment and Integration - LLaMA/Vicuna/Bloom 7B 06/2023
AntGPT [166] 4 CLIP ViT-L/14 - A6000 GPUs Llama2 7B 07/2023
MovieChat [241] 2048 ViT-G/14, EVA-CLIP Q-former+LSTM - GPT-3.5/Claude - 07/2023
FAVOR [282] Varying InstructBLIP ViT-G/14 Causal Q-Former - Vicuna 7B/13B 10/2023
Auto-AD II [275] Varying CLIP ViT-B/32 - - GPT-2 - 10/2023
LLaMA-VQA [240] - CLIP ViT-L/14 Linear Layer 8 A6000 GPUs LLaMA 7B 10/2023
MM-VID [266] Varying - - - GPT4 - 10/2023
Video-LLaVA [239] 8-32 LanguageBind MLP Projection Layer - Vicuna/LLaVA 7B/13B 11/2023
@@
PG-Video-LLaVA [277] Varying CLIP ViT-L/14 MLP Projection Layer 4 A100 GPUs Vicuna/LLaVA 7B/13B 11/2023
SeViLA [255] 32 BLIP-2 ViT Q-former 4 48GB A6000 GPUs MiniGPT4 - 11/2023
@@
Vamos [274] Varying CLIP ViT-L/14, BLIP-2 ViT LORA / LLaMA-Adapter A6000 GPU LLaMA/LLaMA2 7B 11/2023
VTimeLLM [258] 100 CLIP ViT-L/14 Linear Layer 1 RTX-4090 GPU Vicuna 7B/13B 11/2023
Chat-UniVi [237] 64 ViT-L/14, ViT-G Spatial Temporal Merging - Vicuna 7B 11/2023
MM-narrator [273] Varying CLIP ViT - - GPT-4 - 11/2023
@@
GPT4Video [265] Varying CLIP ViT-L/14 Cross-Attention 8 A100 GPUs LLaMA 7B 11/2023
LLaMA-VID [238] 1fps CLIP ViT-L/14 Projector 8 A100 GPUs Vicuna 7B/13B 11/2023
LLoVi [163] 0.5fps CLIP, TimeSformer, BLIP - 32 RTX 3090 GPUs GPT-4 - 12/2023
AV-LLM [231] 32 CLIP ViT-L/14 Linear Projectors 8×A100 GPUs Vicuna 7B 12/2023
Emu2 [232] Varying EVA-02-CLIP-E-plus Linear Projection - Emu2-Chat 38B 12/2023
Grounding-Prompter [164] Varying BLIP - - GPT-3.5-turbo-16k - 12/2023
VIDOSC [165] 1fps CLIP - - GPT4 - 12/2023
LifelongMemory [186] 360 CLIP ViT-L/14, TimeSformer - - GPT-4 - 12/2023
Merlin [269] 8 CLIP ViT-L/14 2D Conv + Linear Layer - Vicuna v1.5 7B 12/2023
@@
MMICT [233] 16 BLIP-2 ViT-G/14 Transformers + Linear Projection 4 V100 GPUs FlanT5XL, OPT 2.7B 12/2023
OneLLM [264] - CLIP ViT-L/14 K Transformer Projectors + 1 MLP 16 A100 GPUs LLaMA2 7B 12/2023
SUM-shot [267] 4 EVA-CLIP ViT-G/14 Q-former 16 A100-80G GPUs Vicuna-v0 7B 12/2023
@@
TimeChat [257] 96 BLIP-2 ViT-G/14 Video Q-Former+Linear Layer+LoRA - LLaMA-2 7B 12/2023
@@
VaQuitA [234] 100 CLIP ViT-L/14 Video Perceiver + VQ-Former 8 A100 80GB GPUs Llama2 7B 12/2023
VILA [235] 8 CLIP ViT-L/14 Linear Projector 16 A100 GPUs Llama-2 7B/13B 12/2023
@@
Vista-LLaMA [236] - EVA-CLIP, CLIP ViT-L/14 Sequential Q-Former with linear layer 8 A100 80GB GPUs Vicuna 7B 12/2023
ProViQ [187] - TimeSformer - one A100 GPU gpt-3.5-turbo - 12/2023
@@
COSMO [219] 3 CLIP ViT-L/14 Gated Cross-Attention Layers 128 V100 GPUs OPT-IML/RedPajama/Mistral 1.8/3/7B 01/2024
VidDetours [263] - InternVideo Linear Layer 8 A100 GPUs LLaMA-2 13B 01/2024
DoraemonGPT [184] - - - - GPT-3.5-turbo - 01/2024
GroundingGPT [256] 64 CLIP ViT-L/14 Q-former, MLP 8 A100 GPUs Vicuna-v1.5 7B 01/2024
GCG [209] 4 EVA-CLIP Q-former - IntructBLIP-Vicuna 7B 01/2024
IVAwithLLM [226] - no mention Linear Layer 8 A100 GPUs Vicuna 7B 02/2024
LSTP [227] Varying BLIP-2 ViT-G/14 Q-former 1 A100 GPU Vicuna 7B 02/2024
@@
LVCHAT [228] Varying UMT-L QFormer + Linear Layer 4 A6000 GPUs Vicuna 7B 02/2024
Momentor [262] 300 CLIP ViT-L/14 Projection Layer 8 A1000 GPUs LLaMA 7B 02/2024
OSCaR [229] - CLIP ViT-L/14 - - Vicuna 7/13B 02/2024
Slot-VLM [230] 1 CLIP ViT-L/14 Linear Layer 1 A100 GPU Vicuna 7B 02/2024
Video ReCap [162] 4 TimeSformer - 8 V100 GPUs GPT-2 1.5B 02/2024
MVU [161] 16 - - 2 A5000 GPUs Mistral/LLaMA-2/Gemma 7B/13B 03/2024
VideoAgent [180] Varying ViCLIP - - GPT-4 - 03/2024
VideoAgent [181] 1 FPS EVA-CLIP-8B-plus - 1 A6000 GPU GPT-4 - 03/2024
VTG-GP [261] 0.5 FPS EVA-CLIP - 8 RTX-3090 GPUs LLaMA-2 7B 03/2024
VURF [182] - - - - GPT-3.5 - 03/2024
KEPP [183] - - - 2 A100 GPUs LLaMA-2 13B/70B 03/2024
GPTSee [172] - CLIP ViT-B/32, SlowFast Cross-attention 8 RTX 3090 GPUs - - 03/2024
HawkEye [252] Varying UMT-L Q-former + LORA 8 V100 GPUs GPT-4 - 03/2024

### v8

@@ II-B Background for LLMs
  • Scaling Laws [kaplan2020scaling]: With a significant extension of the model size (number of parameters), the pertaining data size, and computational resources, the performance of the model exhibits a pattern of regular growth, which can help researchers and engineers predict performance improvements or make effective decisions on model design and training.
  • Emergent Abilities [zhao2023survey]: When the parameter size and volume of training data for a large language model exceed a certain magnitude, some novel capabilities emerge, such as in-context learning, instruction following, and step-by-step reasoning. In-context learning allows the model to learn and make predictions based on the context provided within the input text without requiring explicit retraining. Instruction following enables the model to perform tasks based on natural language instructions. Step-by-step reasoning, e.g. chain-of-thought (CoT), allows the model to follow a logical sequence of steps to arrive at a conclusion, which is particularly useful for solving complex problems.
LLMs possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodal Large Language Models (MLLMs) [zhu2023minigpt, liu2023visual, chen2023shikra, xuan2023pink, hua2024finematch] typically incorporate multimodal encoders, cross-modality aligners, and an LLM core structure. By combining multimodal encoders with the LLM, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Summarizer: In this subcategory, the primary function of the LLM is to summarize the analysis obtained from the Video Analyzers. The summarization approach varies based on the prompts provided to the LLM, ranging from highly condensed summary texts and captions to comprehensive summaries for answering specific questions. Notably, in Video Analyzer $\times$ LLM as Summarizer systems, the information flow is usually unidirectional (see Figure 4), with data flowing from the video to the Video Analyzer and then to the LLM, without any reverse process. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Summarizer category include: LaViLa [zhao2023lavila], VLog [VLogGithub], VAST [chen2023vast], AntGPT [zhao2023antgpt], VIDOSC [Xue2024VIDOSC], Grounding-Prompter [Chen2023Grounding-prompter], LLoVi [Zhang2024LLoVi], Video ReCap [Islam2024VideoReCap], MVU [Ranasinghe2024MVU], LangRepo [Kahatapitiya2024LangRepo], IG-VLM [Kim2024IG-VLM], MoReVQA [min2024morevqa], MM-Screenplayer [wu2024mmscreenplayer], GIT-LLaVA [kalarani2024gitllava], etc.
  • LLM as Manager: In this subcategory, the LLM primarily coordinates the overall system’s operation. It may actively generate commands to invoke various Video Analyzers to produce the desired results for the user, call the Video Analyzer and further process the obtained analysis before returning it to the user, or engage in multiple rounds of interaction with the Video Analyzer. Compared to the LLM as Summarizer category, the LLM as Manager category is more flexible and can be distinguished by its information flow complexity. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Manager category include: ViperGPT [suris2023vipergpt], Video ChatCaptioner [chen2023videochatcap], ChatVideo [wang2023chatvideo], AssistGPT [gao2023assistgpt], HuggingGPT [shen2024hugginggpt], Hawk [tang2024hawk], ProViQ [Choudhury2023ProViQ], LifelongMemory [Wang2024LifelongMemory], DoraemonGPT [Yang2024DoraemonGPT], KEPP [Nagasinghe2024KEPP], VURF [Mahmood2024VURF], VideoAgent (by Stanford) [Wang2024VideoAgent], VideoAgent (by PKU) [Fan2024VideoAgent], TV-TREES [S2024TV-TREES], SCHEMA [Niu2024SCHEMA], RAVA [Cao2024RAVA], GPTSee [Sun2024GPTSee], TraveLER [shang2024traveler], LAVAD [zanella2024lavad], VideoTree [wang2024videotree], LVNet [park2024lvnet], OmAgent [zhang2024omagent], DrVideo [ma2024drvideo], etc.
@@ III-A2 Video Embedder $\times$ LLM
A Video Embedder typically refers to a visual backbone/video encoder, such as ViT or CLIP, used to convert input videos into vector representations, known as video embeddings or video tokens. Some Embedders encode other modalities within the video, such as audio (e.g., CLAP [elizalde2023clap]), which are also categorized under Video Embedder here (note that we do not consider the LLM’s tokenizer as an embedder). Unlike the text generated by the Video Analyzer, the vectors generated by the Video Embedder cannot be directly utilized by the LLM and usually require an adapter to map these embeddings from the vision’s (or other modalities’) semantic space to the text semantic space of the LLM’s input tokens.
@@
  • LLM as Text Decoder: In this subcategory, the LLM receives embeddings from the Video Embedder as input and decodes them into text outputs based on prompts or instructions. These tasks generally do not require fine-grained understanding or precise spatiotemporal localization, focusing mainly on general QA or captioning. Thus, the Vid-LLM behaves like a standard LLM during decoding. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Text Decoder category include: VideoLLM [chen2023videollm], Otter [li2023otter], Video-LLaMA [zhang2023video], Video-ChatGPT [maaz2023video], VALLEY [luo2023valley], Macaw-LLM [lyu2023macaw], MovieChat [song2023moviechat], Video-LLaVA [lin2023video], Chat-UniVi [Jin2024Chat-UniVi], Vista-LLaMA [Ma2023Vista-LLaMA], VILA [Lin2024VILA], GPT4Video [wang2023gpt4video], MovieLLM [Song2024MovieLLM], InternVideo2 [Wang2024InternVideo2], MiniGPT4-Video [ataallah2024minigpt4video], VideoChat2 [li2024mvbench], VideoLLaMA 2 [cheng2024videollama2], etc. See Figure 5 for the complete list.
  • LLM as Regressor: In this subcategory, the LLM receives embeddings from the Video Embedder as input and, like the Text Decoder, can output text. However, unlike the Text Decoder, the LLM as Regressor can also predict continuous values, such as timestamp localization in videos and bounding box coordinates for object trajectories, functioning similarly to a regressor performing regression tasks, even though it is fundamentally performing classification. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Regressor category include: VTimeLLM [huang2023vtimellm], SeViLA [Yu2023SeViLA], TimeChat [Ren2024TimeChat], GroundingGPT [Li2024GroundingGPT], OmniViD [Wang2024OmniViD], LITA [Huang2024LITA], HawkEye [Wang2024HawkEye], Elysium [wang2024elysium], AVicuna [tang2024avicuna], V2Xum-LLaMA [hua2024v2xum], VLM4HOI [bansal2024hoiref], VideoLLM-online [chen2024videollmonline], Holmes-VAD [zhang2024holmesvad], etc.
  • LLM as Hidden Layer: In this subcategory, the LLM also receives video embeddings as input but does not directly output text. Instead, it connects to a specially designed task-specific head to perform actual regression tasks, such as event time localization or object bounding box prediction in videos, while maintaining the LLM’s text output capability. Examples of Vid-LLMs in the Video Embedder $\times$ LLM as Hidden Layer category include: GPT4Video [wang2023gpt4video], OneLLM [Han2023OneLLM], VidDetours [Ashutosh2024VidDetours], Momentor [Qian2024Momentor], VTG-GPT [Xu2024VTG-GPT], VITRON [fei2024vitron], VTG-LLM [guo2024vtgllm], etc.
@@
This category of Vid-LLMs is relatively rare. As the name suggests, it involves simultaneously using a Video Analyzer to obtain textual analysis of the video and a Video Embedder to encode the video into embeddings. The LLM receives both types of inputs along with other prompts/instructions and outputs responses to complete tasks. The subcategories here can flexibly be any of the Summarizer/Manager/Text Decoder/Regressor/Hidden Layer categories. Vid-LLMs in the (Analyzer + Embedder) $\times$ LLM category include: Vid2Seq [yang2023vid2seq], VideoChat [li2023videochat], MM-VID [lin2023mmvid], Auto-AD II [han2023autoad], Vamos [Wang2024Vamos], PG-Video-LLaVA [munasinghe2023pg], MM-Narrator [zhang2023mm-narrator], SUM-shot [Han2023SUM-shot], Merlin [yu2023merlinempowering], Uni-AD [Wang2024Uni-AD], Vriptor [yang2024vript], etc.
@@ III-B1 Training-free Vid-LLMs
Many Vid-LLMs systems are built on powerful LLMs with strong zero-shot, in-context learning, and Chain-of-Thought capabilities. These systems do not require training in the parameters of the LLM or other modules. Most Vid-LLMs in the Video Analyzer $\times$ LLM category are training-free because the information from the video and other accompanying modalities has already been parsed into text. At this point, the video understanding task has been transformed into a text understanding task. Since LLMs can unify almost all NLP tasks into generation tasks, they can also handle many video understanding tasks. SlowFast-LLaVA [xu2024slowfastllava] is a training-free Video LLM that uses a two-stream input design to capture both spatial semantics and temporal context without fine-tuning and demonstrates capabilities across various video understanding benchmarks.
@@
Socratic Models [zeng2022socratic] Varying CLIP ViT-L/14 - 1 V100 GPU RoBERTa, GPT-3 - 05/2022
LaViLa [zhao2023lavila] 4 TimeSformer-B/L Cross-Attention 32 V100 GPUs GPT-2 XL frozeen - 12/2022
Vid2Seq [yang2023vid2seq] 100 CLIP ViT-L/14 Transformer Encoder 64 TPU v4 T5 0.2B 02/2023
ViperGPT [suris2023vipergpt] - - - - - - 03/2023
Vid ChatCaptioner [chen2023videochatcap] 100 - BLIP-2 24G GPU ChatGPT 20B 04/2023
VLog [VLogGithub] - CLIP ViT-G - - ChatGPT 20B 04/2023
ChatVideo [wang2023chatvideo] - - - - ChatGPT 20B 04/2023
VideoChat [li2023videochat] 4-32 ViT-G MLP+Q-fomer 1 A10 GPU StableVicuna 7B 05/2023
ChatBridge [zhao2023chatbridge] 4 ViT-G Perceiver 8 A100 GPUs GPT-4 - 05/2023
VAST [chen2023vast] 4,8 EVA-CLIP, BEATs, BERT-B - 64 V100 GPUs Vicuna 13B 05/2023
Otter [li2023otter] 4-8 CLIP ViT-L/14 - 4 RTX-3090 GPUs LLaMA 7B 05/2023
@@
VideoLLM [chen2023videollm] Varying 7 Task-Specific Video Encoders Linear Layer - GPT-2/T5/OPT/LLaMA 1.5/6.5/6.7/7B 05/2023
AssistGPT [gao2023assistgpt] 1/3 FPS - - 4 A5000 GPUs Vicuna 7B 06/2023
Video-LLaMA [zhang2023video] 8 CLIP ViT-G Video Q-former - Vicuna 7B 06/2023
@@
Video-ChatGPT [maaz2023video] 100 CLIP ViT-L/14 Linear Layer 8 A100 GPUs Vicuna-v1.1 7B 06/2023
LLMVA-GEBC [tang2023llmva] 96 CLIP ViT-G Video Q-former 2 A6000 GPUs OPT 13B 06/2023
mPLUG-video [xu2023youku] 8 TimeSformer - - GPT/Blood - 06/2023
VALLEY [luo2023valley] Varying CLIP ViT-L/14 Projection Layer 8 A100 GPUs StableVicuna 7B/13B 06/2023
@@
Macaw-LLM [lyu2023macaw] Varying CLIP VIT-B/16 Alignment and Integration - LLaMA/Vicuna/Bloom 7B 06/2023
AntGPT [zhao2023antgpt] 4 CLIP ViT-L/14 - A6000 GPUs Llama2 7B 07/2023
MovieChat [song2023moviechat] 2048 ViT-G/14, EVA-CLIP Q-former+LSTM - GPT-3.5/Claude - 07/2023
FAVOR [sun2023finegrained] Varying InstructBLIP ViT-G/14 Causal Q-Former - Vicuna 7B/13B 10/2023
Auto-AD II [han2023autoad] Varying CLIP ViT-B/32 - - GPT-2 - 10/2023
LLaMA-VQA [ko2023llamavqa] - CLIP ViT-L/14 Linear Layer 8 A6000 GPUs LLaMA 7B 10/2023
MM-VID [lin2023mmvid] Varying - - - GPT4 - 10/2023
Video-LLaVA [lin2023video] 8-32 LanguageBind MLP Projection Layer - Vicuna/LLaVA 7B/13B 11/2023
@@
PG-Video-LLaVA [munasinghe2023pg] Varying CLIP ViT-L/14 MLP Projection Layer 4 A100 GPUs Vicuna/LLaVA 7B/13B 11/2023
SeViLA [Yu2023SeViLA] 32 BLIP-2 ViT Q-former 4 48GB A6000 GPUs MiniGPT4 - 11/2023
@@
Vamos [Wang2024Vamos] Varying CLIP ViT-L/14, BLIP-2 ViT LORA / LLaMA-Adapter A6000 GPU LLaMA/LLaMA2 7B 11/2023
VTimeLLM [huang2023vtimellm] 100 CLIP ViT-L/14 Linear Layer 1 RTX-4090 GPU Vicuna 7B/13B 11/2023
Chat-UniVi [Jin2024Chat-UniVi] 64 ViT-L/14, ViT-G Spatial Temporal Merging - Vicuna 7B 11/2023
MM-narrator [zhang2023mm-narrator] Varying CLIP ViT - - GPT-4 - 11/2023
@@
GPT4Video [wang2023gpt4video] Varying CLIP ViT-L/14 Cross-Attention 8 A100 GPUs LLaMA 7B 11/2023
LLaMA-VID [li2023llama-vid] 1fps CLIP ViT-L/14 Projector 8 A100 GPUs Vicuna 7B/13B 11/2023
LLoVi [Zhang2024LLoVi] 0.5fps CLIP, TimeSformer, BLIP - 32 RTX 3090 GPUs GPT-4 - 12/2023
AV-LLM [Shu2023AV-LLM] 32 CLIP ViT-L/14 Linear Projectors 8×A100 GPUs Vicuna 7B 12/2023
Emu2 [Sun2024Emu2] Varying EVA-02-CLIP-E-plus Linear Projection - Emu2-Chat 38B 12/2023
Grounding-Prompter [Chen2023Grounding-prompter] Varying BLIP - - GPT-3.5-turbo-16k - 12/2023
VIDOSC [Xue2024VIDOSC] 1fps CLIP - - GPT4 - 12/2023
LifelongMemory [Wang2024LifelongMemory] 360 CLIP ViT-L/14, TimeSformer - - GPT-4 - 12/2023
Merlin [yu2023merlinempowering] 8 CLIP ViT-L/14 2D Conv + Linear Layer - Vicuna v1.5 7B 12/2023
@@
MMICT [Chen2023MMICT] 16 BLIP-2 ViT-G/14 Transformers + Linear Projection 4 V100 GPUs FlanT5XL, OPT 2.7B 12/2023
OneLLM [Han2023OneLLM] - CLIP ViT-L/14 K Transformer Projectors + 1 MLP 16 A100 GPUs LLaMA2 7B 12/2023
SUM-shot [Han2023SUM-shot] 4 EVA-CLIP ViT-G/14 Q-former 16 A100-80G GPUs Vicuna-v0 7B 12/2023
@@
TimeChat [Ren2024TimeChat] 96 BLIP-2 ViT-G/14 Video Q-Former+Linear Layer+LoRA - LLaMA-2 7B 12/2023
@@
VaQuitA [Wang2023VaQuitA] 100 CLIP ViT-L/14 Video Perceiver + VQ-Former 8 A100 80GB GPUs Llama2 7B 12/2023
VILA [Lin2024VILA] 8 CLIP ViT-L/14 Linear Projector 16 A100 GPUs Llama-2 7B/13B 12/2023
@@
Vista-LLaMA [Ma2023Vista-LLaMA] - EVA-CLIP, CLIP ViT-L/14 Sequential Q-Former with linear layer 8 A100 80GB GPUs Vicuna 7B 12/2023
ProViQ [Choudhury2023ProViQ] - TimeSformer - one A100 GPU gpt-3.5-turbo - 12/2023
@@
COSMO [Wang2024COSMO] 3 CLIP ViT-L/14 Gated Cross-Attention Layers 128 V100 GPUs OPT-IML/RedPajama/Mistral 1.8/3/7B 01/2024
VidDetours [Ashutosh2024VidDetours] - InternVideo Linear Layer 8 A100 GPUs LLaMA-2 13B 01/2024
DoraemonGPT [Yang2024DoraemonGPT] - - - - GPT-3.5-turbo - 01/2024
GroundingGPT [Li2024GroundingGPT] 64 CLIP ViT-L/14 Q-former, MLP 8 A100 GPUs Vicuna-v1.5 7B 01/2024
GCG [Wang2024GCG] 4 EVA-CLIP Q-former - IntructBLIP-Vicuna 7B 01/2024
IVAwithLLM [Li2024IVAwithLLM] - no mention Linear Layer 8 A100 GPUs Vicuna 7B 02/2024
LSTP [Wang2024LSTP] Varying BLIP-2 ViT-G/14 Q-former 1 A100 GPU Vicuna 7B 02/2024
@@
LVCHAT [Wang2024LVCHAT] Varying UMT-L QFormer + Linear Layer 4 A6000 GPUs Vicuna 7B 02/2024
Momentor [Qian2024Momentor] 300 CLIP ViT-L/14 Projection Layer 8 A1000 GPUs LLaMA 7B 02/2024
OSCaR [Nguyen2024OSCaR] - CLIP ViT-L/14 - - Vicuna 7/13B 02/2024
Slot-VLM [Xu2024Slot-VLM] 1 CLIP ViT-L/14 Linear Layer 1 A100 GPU Vicuna 7B 02/2024
Video ReCap [Islam2024VideoReCap] 4 TimeSformer - 8 V100 GPUs GPT-2 1.5B 02/2024
MVU [Ranasinghe2024MVU] 16 - - 2 A5000 GPUs Mistral/LLaMA-2/Gemma 7B/13B 03/2024
VideoAgent [Fan2024VideoAgent] Varying ViCLIP - - GPT-4 - 03/2024
VideoAgent [Wang2024VideoAgent] 1 FPS EVA-CLIP-8B-plus - 1 A6000 GPU GPT-4 - 03/2024
VTG-GP [Xu2024VTG-GPT] 0.5 FPS EVA-CLIP - 8 RTX-3090 GPUs LLaMA-2 7B 03/2024
VURF [Mahmood2024VURF] - - - - GPT-3.5 - 03/2024
KEPP [Nagasinghe2024KEPP] - - - 2 A100 GPUs LLaMA-2 13B/70B 03/2024
GPTSee [Sun2024GPTSee] - CLIP ViT-B/32, SlowFast Cross-attention 8 RTX 3090 GPUs - - 03/2024
HawkEye [Wang2024HawkEye] Varying UMT-L Q-former + LORA 8 V100 GPUs GPT-4 - 03/2024